OpenStudy (anonymous):

What is the slope of the line 6x + 8y = 12?

OpenStudy (anonymous):

Do you know the equation to find slope?

OpenStudy (zale101):

You want to set your equation to y=mx+b by that, subtract 6x to both sides to leave y by itself, then divide each term by 8. Do these and let me know.

OpenStudy (anonymous):

6x + 8y = 12 8y = -6x + 12 y = -6/8x + 12/8 y = -3/4x + 3/2 The slope is -3/4

OpenStudy (anonymous):

Thank you!